OpenStudy (cutiekawdfi):

Write an equation of an ellipse centered at the origin,satisfying the given conditions. 1. vertex (0, square root 29) ; co vertex (-5,0)

OpenStudy (cutiekawdfi):

@Tommynaiter

OpenStudy (anonymous):

An ellipse has the genera equation: \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]Where: x,y are the coordinates of any point on the ellipse, a, b are the radius on the x and y axes

OpenStudy (cutiekawdfi):

im confused

OpenStudy (cutiekawdfi):

@Tommynaiter